Planet X has an estimated mass of 1.78x1027 kg, while its diameter is approximately 142,984 km. Consider the planet as a perfect sphere. What is the density of Planet X in g/cm3 ? Volsphere = 4 /3 πr 3
Answer #1

Answer

1.16g/cm3

Explanation

Radius = diameter/2

Radius of the planet X = 142984km/2 = 71492km = 71492000m = 7149200000cm

Volume of the planet X = (4/3) × 3.143 × ( 7149200000cm)3 = 1.53× 1030cm3

mass of the planet X = 1.78 × 1027 kg = 1.78 × 103​​​​​0g

Density = mass /volume

Density of the planet X = 1.78 ×1030g/1.53×1030cm3 = 1.16g/cm3